# Application of Optimal Control to Time-Resolution Protocol for Quantum Sensing

### v1

@@ Abstract
Time-resolution protocol of quantum sensing aims to measure the fast temporal variation of an external field and demands a high field sensitivity in a short interrogation time $\tau$. Since any operation that evolves the quantum state takes time and is counted as part of the interrogation, evaluating the performance of time-resolution protocol requires a complete end-to-end description of the measurement process. In particular, the initial state has to be one of the sensor qubit’s eigenstates in the absence of external fields, and the final projective measurements must be performed in the same eigenstate basis. Building upon prior works which proposed limits for time-resolved sensing using a quantum sensor, we apply optimal control theory to optimize the time-resolution protocol. Our analysis indicates that there exists a critical interrogation time $T^{*}$: when $\tau<T^{*}$ the optimal protocol is purely bang-bang; when $\tau>T^{*}$ the optimal protocol involves a singular control during the interrogation. In the short-$\tau$ regime, which is relevant to high time resolution, we propose a “detune protocol” that involves only smooth control during the entire interrogation. As the discontinuities of control pose the main obstacles to experimental realization, we expect the presented detune protocol to be practically useful. In the long-$\tau$ regime, the optimal protocol closely resembles the Ramsey sequence; protocols based on maximizing Quantum Fisher Information are constructed to highlight the difference between the theoretically optimal and practically implementable measurements. Effective use of the time-resolution protocol requires a setup where the unknown time-domain signal of interest can be identically and repeatedly generated. As a potentially relevant application, we outline the calibration of baseband flux pulse distortion in the control of superconducting qubits.
@@ I Introduction
Quantum sensing generally refers to the technology that explicitly makes use of quantum phenomena to achieve highly sensitive and precise measurement Degen et al. (2017). Notable examples include employing the squeezed state of light to enhance the sensitivity of laser interferometers Xiao et al. (1987); Grangier et al. (1987); Breitenbach et al. (1997); Tse et al. (2019); Acernese et al. (2019), using Nitrogen-Vacancy (NV) centers in diamond to measure signals generated by sub-micron objects Barry et al. (2020); Glenn et al. (2018); Ku et al. (2020); Turner et al. (2020), and utilizing the multi-level interference spectroscopy of Rydberg atoms to determine the amplitude of microwave and terahertz electromagnetic signals Saffman et al. (2010); Sedlacek et al. (2012, 2013); Fan et al. (2014); Meyer et al. (2018); Holloway et al. (2019); Downes et al. (2020); Cloutman et al. (2024). Intrinsic quantum advantage amounts to identifying and preparing the quantum state that is most sensitive to the field to be measured. The sensing performance can be quantified by the Quantum Fisher Information (QFI) Helstrom (1976); Holevo (2011); given an unbiased estimator, the inverse of QFI gives the lower bound (Cramér-Rao bound) of the variance of the parameter to be estimated. QFI, which depends solely on the quantum state and not on the specific measurement observables, is very valuable in both formal analysis and numerical simulations. From a theoretical point of view, QFI provides a framework for comparing performances of different states. Given a sensing setup, analysis of QFI identifies its theoretical bound (known as the Heisenberg limit) and the corresponding best-sensing state (equal superposition of eigenstates with the largest and smallest eigenvalues) Giovannetti et al. (2006, 2011). From a practical perspective, maximizing QFI over the available control parameters offers a concrete means – or at least a guide – for approaching the quantum limit Liu and Yuan (2017a, b); Haine and Hope (2020); Lin et al. (2021, 2022).
@@
For realistic implementations, both the initial state and the measurement observable have to be specified to determine the overall sensing performance; they are especially essential for the time-resolution protocols. The objective of the time-resolution protocol is to maximize the sensitivity with respect to the to-be-measured field in a short interrogation time $\tau$. Because any operation that evolves the quantum state takes time, the description of the time-resolution protocol should begin with the qubit in one of its eigenstates in the absence of external fields, and end up with a projective measurement in the same eigenstate basis. In Ref. Herb and Degen (2024), Herb and Degen propose a time-resolution protocol based on the Ramsey sequence Ramsey (1950) and the Rotating Wave Approximation (RWA). Their protocol, which will be served as the reference for comparison, consists of two sequential pulses of equal duration, one corresponding to $\hat{\sigma}_{y}$ and the other to $\hat{\sigma}_{x}$ in the rotating frame, and has been experimentally demonstrated Herb et al. (2025).
@@
In this work, we apply Optimal Control Theory (OCT) to analyze the time-resolution protocol. OCT provides an efficient way to minimize a user-defined terminal cost subject to the dynamics that contains a time-dependent scalar control Luenberger (1979); Liberzon (2012); Heinz Schattler (2012); Pontryagin (1987). It has emerged as a valuable tool for engineering a wide range of quantum systems Rembold et al. (2020); Boscain et al. (2021); Magann et al. (2021); Ansel et al. (2024), such as stabilizing ultracold molecules Koch et al. (2004), optimizing the performance in nuclear magnetic resonance measurement Lapert et al. (2014); Kobzar et al. (2012); Dridi et al. (2020), and cooling of quantum systems Stefanatos et al. (2010, 2011); Rahmani et al. (2013). Apart from the better sensitivity, which is expected as the optimal control gives the best possible performance, the general behavior of the optimal protocols is qualitatively similar to that of the reference protocols. In the short-$\tau$ regime which is relevant to high time resolution, we propose a detune protocol that only involves the smooth control field over the entire interrogation and performs comparably to the reference protocol. Given that discontinuities in the control field pose the primary experimental challenge, we expect the detune protocol to be practically useful. One bottleneck of meaningfully utilizing the time-resolution protocol is the need to generate the identical, yet unknown time-domain signal repeatably. We suggest that it can be used to calibrate the pulse distortion in fast flux control in superconducting qubits Rol et al. (2020); Hellings et al. (2025); Li et al. (2025), which is an essential but cumbersome process for calibrating quantum gates, and is increasingly detrimental as the number of qubits gets larger. We discuss the potential to substantially reduce the total calibration time by leveraging our protocol to perform sequential measurements of the same distorted pulse. The rest of the paper is organized as follows. Section II defines the problem of time resolution and provides the necessary background. Section III presents our main results, including the general behavior from optimal control and a detailed description of the detune protocol. Section IV discusses practical aspects, including an outline of a realistic application. Section V provides a brief conclusion.
@@ II.1 Overview and problem statement
In this subsection we formulate the design of the time-resolution protocol for quantum sensing as an optimal control problem. The sensor qubit is described by the Hamiltonian Herb and Degen (2024)
@@
where $\omega_{0}$ is the natural angular frequency of the sensor qubit, $\delta\omega$ is the field which we want to measure, and $u(t)$ is a scalar field which we can control. We shall take $\hbar\equiv 1$, $\omega_{0}\equiv 1$ in our simulations but explicitly keep $\omega_{0}$ in all expressions; within this convention energy is measured in units of $\hbar\omega_{0}$ and time $\omega_{0}^{-1}$. $|0\rangle$ and $|1\rangle$ are eigenstates of $\hat{\sigma}_{z}$ with eigenvalues of 1 and -1 respectively. They are eigenstates without external fields (i.e., $\delta\omega=u=0$) and will be referred to as the natural eigenstates of the sensor qubit.
The performance of the time-resolution protocol is quantified by its measurement sensitivity with respect to $\delta\omega$ given an (short) interrogation time $\tau$. Because any operation contributes to the interrogation time, an end-to-end description of the measurement process is required for performance evaluation. In particular the initial state and the basis of the final projection measurement have to be the sensor qubit’s natural eigenstates. Without loss of generality we choose the initial state to be $|0\rangle$ and measure the probability that ends up with the same $|0\rangle$ state; the outcome probability can be expressed as
@@
$p_{0}$ is the probability with zero field (i.e., $\delta\omega=0$) and is independent of $\delta\omega$; the linear coefficient of $\delta\omega$, denoted as $\eta$, is the “measurement sensitivity” Herb and Degen (2024):
@@ II.2 Reference: resonant YX protocol
The reference protocol, which will be termed as the resonant YX protocol or simply the YX protocol, is the one proposed by Herb and Degen in Ref. Herb and Degen (2024); its control waveform is
@@
where $t_{\text{QSL}}=\frac{\pi}{u_{\text{max}}}$ is referred to as the quantum speed limit Herb and Degen (2024); Qua ; Mandelstam and Tamm (1945); Margolus and Levitin (1998); Deffner and Campbell (2017). The rationale behind the protocol naming is as follows: the “resonant” indicates that the applied frequency is identical to the natural frequency; “YX” that in RWA, $\cos(\omega_{0}t+\frac{\pi}{2})\hat{\sigma}_{x}$ and $\cos(\omega_{0}t)\hat{\sigma}_{x}$ respectively lead to $\hat{\sigma}_{y}$ (Y operation) and $\hat{\sigma}_{x}$ (X operation) in the rotating frame. Overall the YX protocol identifies a critical interrogation time $t_{\text{QSL}}$, below which the protocol is composed of equal-duration Y and X operations, and above which the protocol involves a free evolution between Y and X operations.
@@
Here $K$ is a protocol dependent convolution kernel; for constant $\delta\omega$ one recovers $\eta=\int_{-\tau/2}^{\tau/2}\text{d}tK(t)$. In Appendix A (see also Supplementary of Ref. Herb and Degen (2024)) we provide a simple numerical method to construct the kernel of Eq. (8). Once $K$ is known, the time resolution shorter than $\tau$ is in principle possible Herb et al. (2025).
@@ II.3 Optimal control
In this subsection we summarize the important functions introduced in OCT. Numerous comprehensive reviews have been reported in the literature Boscain et al. (2021); Magann et al. (2021), and its applications specific to the unitary qubit can be found in Ref. Lin et al. (2025). Here, special attention is devoted to the cost function that involves a derivative of the wave function.
@@
OCT is a powerful tool for dynamics-constrained optimization. Given the dynamics and the cost to minimize, it introduces the adjoint variables $|\pi(t)\rangle$, the control-Hamiltonian $\mathcal{H}_{\text{oc}}(t)$, and the switching function $\Phi(t)$ [defined shortly in Eqs. (12)] that are computationally accessible and accurately characterize the system response to some external or model variables. One non-trivial complication from the cost $\mathcal{C}_{\eta^{2}}$ is its dependence on $|\psi_{1}(\tau)\rangle=\frac{\partial}{\partial(\delta\omega)}|\psi_{0}(\tau)\rangle$. To cope with this additional dependence we augment the system by regarding $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$ as independent variables; more details can be found in Ref. Lin et al. (2021, 2020).
@@
We also consider maximizing QFI Lin et al. (2021) where the terminal cost is negative QFI:
@@
Eqs. (12) plus Eq. (13) [or Eq. (14)] completely determine quantities introduced by OCT. They are used to express two general optimality conditions con . First, the optimal control is
@@
We emphasize that although the physical system considered here is a qubit with unitary dynamics [i.e., Eq. (1)], the augmented dynamics specified in Eq. (12c) is neither planar (i.e., two real-valued variables) nor unitary. For these reasons certain properties specific to the unitary qubit given in Ref. Lin et al. (2025) do not hold anymore. Particularly we point out that the BB control does not have identical bang duration and the singular control does not require $u^{\text{sing}}=0$.
@@ II.4 Practical use of OCT functions
We conclude this section by highlighting the usefulness of OCT. The most practically useful quantity is undoubtedly the switching function $\Phi$. By sequentially solving the dynamics twice, one forward in time for $|\psi\rangle$ and one backward $|\pi\rangle$, one is able to obtain the gradient of $\mathcal{C}$ with respect to all $u(t_{i})$ Khaneja et al. (2005). Three applications used in the work are highlighted. First $\Phi$ is primarily used to construct a numerical $u^{*}(t)$ by some gradient-based (or quasi-Newton) algorithm that involves
@@
For the free-form parametrization one divides $\tau$ into $N_{t}$ intervals and assumes a piecewise constant waveform. In our simulations $N_{t}$ ranges from 100 to 2000 (depending on $\tau$); we ensure that the result has a negligible change upon increasing $N_{t}$’s and that $N_{t}$ is sufficiently large so that all optimality conditions are numerically satisfied Lin et al. (2021). Empirically we find that when $u^{*}(t)$ contains both B and S controls, the convergence of S control is slow due to its vanishing gradient. In this case we can compute Eq. (16) at $n$th iteration, and replace the non-bang controls by the singular values:
@@
Two practical utilizations of control-Hamiltonian $\mathcal{H}_{\text{oc}}(t)\sim\frac{\partial\mathcal{C}}{\partial\tau}$ are outlined. First, the deviation from a constant $\mathcal{H}_{\text{oc}}(t)$ can be used to quantify the solution quality Lin et al. (2021). Second, in a damped quantum system $\mathcal{H}_{\text{oc}}=0$ offers a rigorous criterion for defining the optimal operation time for a quantum task Lin et al. (2022).

### v2

@@ Abstract
Time-resolution protocol of quantum sensing aims to measure the fast temporal variation of an external field and demands a high field sensitivity in a short interrogation time $\tau$. Since any operation that evolves the quantum state takes time and is counted as part of the interrogation, evaluating the performance of time-resolution protocol requires a complete end-to-end description of the measurement process. In particular, the initial state has to be one of the quantum sensor’s eigenstates in the absence of external fields, and the final projective measurements must be performed in the same eigenstate basis. Building upon prior works which proposed limits for time-resolved sensing using a quantum sensor, we apply optimal control theory to optimize the time-resolution protocol. Our analysis indicates that there exists a critical interrogation time $T^{*}$: when $\tau<T^{*}$ the optimal protocol is purely bang-bang; when $\tau>T^{*}$ the optimal protocol involves a singular control during the interrogation. In the short-$\tau$ regime, which is relevant to high time resolution, we propose a “detune protocol” that involves only smooth control during the entire interrogation. As the discontinuities of control pose the main obstacles to experimental realization, we expect the presented detune protocol to be practically useful. In the long-$\tau$ regime, the optimal protocol closely resembles the Ramsey sequence; protocols based on maximizing Quantum Fisher Information are constructed to highlight the difference between the theoretically optimal and practically implementable measurements. Effective use of the time-resolution protocol requires a setup where the unknown time-domain signal of interest can be identically and repeatedly generated. As a potentially relevant application, we outline the calibration of baseband flux pulse distortion in the control of superconducting qubits.
@@ I Introduction
Quantum sensing generally refers to the technology that explicitly makes use of quantum phenomena to achieve highly sensitive and precise measurement RevModPhys.89.035002. Notable examples include employing the squeezed state of light to enhance the sensitivity of laser interferometers PhysRevLett.59.278; PhysRevLett.59.2153; Breitenbach-1997; PhysRevLett.123.231107; PhysRevLett.123.231108, using Nitrogen-Vacancy (NV) centers in diamond to measure signals generated by sub-micron objects RevModPhys.92.015004; Glenn-2018; Ku-2020; PhysRevApplied.14.014097, and utilizing the multi-level interference spectroscopy of Rydberg atoms to determine the amplitude of microwave and terahertz electromagnetic signals RevModPhys.82.2313; Sedlacek-2012; PhysRevLett.111.063001; Fan-2014; 10.1063/1.5028357; 10.1063/1.5099036; PhysRevX.10.011027; PhysRevApplied.21.044025. Intrinsic quantum advantage amounts to identifying and preparing the quantum state that is most sensitive to the field to be measured. The sensing performance can be quantified by the Quantum Fisher Information (QFI) book:Helstrom; book:Holevo; given an unbiased estimator, the inverse of QFI gives the lower bound (Cramér-Rao bound) of the variance of the parameter to be estimated. QFI, which depends solely on the quantum state and not on the specific measurement observables, is very valuable in both formal analysis and numerical simulations. From a theoretical point of view, QFI provides a framework for comparing performances of different states. Given a sensing setup, analysis of QFI identifies its theoretical bound (known as the Heisenberg limit) and the corresponding best-sensing state (equal superposition of eigenstates with the largest and smallest eigenvalues) PhysRevLett.96.010401; AdvancesQuantumMetrology_2011. From a practical perspective, maximizing QFI over the available control parameters offers a concrete means – or at least a guide – for approaching the quantum limit PhysRevA.96.012117; PhysRevA.96.042114; PhysRevLett.124.060402; PhysRevA.103.052607; PhysRevA.105.042621.
@@
For realistic implementations, both the initial state and the measurement observable have to be specified to determine the overall sensing performance; they are especially essential for the time-resolution protocols. The objective of the time-resolution protocol is to maximize the sensitivity with respect to the to-be-measured field in a short interrogation time $\tau$. Because any operation that evolves the quantum state takes time, the description of the time-resolution protocol should begin with the quantum sensor in one of its eigenstates in the absence of external fields, and end up with a projective measurement in the same eigenstate basis. In Ref. PhysRevLett.133.210802, Herb and Degen propose a time-resolution protocol based on the Ramsey sequence PhysRev.78.695 and the Rotating Wave Approximation (RWA). Their protocol, which will be served as the reference for comparison, consists of two sequential pulses of equal duration, one corresponding to $\hat{\sigma}_{y}$ and the other to $\hat{\sigma}_{x}$ in the rotating frame, and has been experimentally demonstrated Herb_2025.
@@
In this work, we apply Optimal Control Theory (OCT) to analyze the time-resolution protocol. OCT provides an efficient way to minimize a user-defined terminal cost subject to the dynamics that contains a time-dependent scalar control book:Luenberger; book:Liberzon; book:GeometricOptimalControl; book:Pontryagin. It has emerged as a valuable tool for engineering a wide range of quantum systems doi:10.1116/5.0006785; PRXQuantum.2.030203; PRXQuantum.2.010101; ansel2024introduction, such as stabilizing ultracold molecules PhysRevA.70.013402, optimizing the performance in nuclear magnetic resonance measurement PhysRevA.90.023411; KOBZAR2012142; PhysRevA.101.012321, cooling of quantum systems PhysRevA.82.063422; doi:10.1137/100818431; PhysRevA.87.043607, and enhancing the yield of the target product in a chemical process PhysRevLett.104.220502; doi:10.1021/acs.jpclett.2c02840; PRXQuantum.5.020303; smith2025interradicalmotionpushmagnetosensing. Apart from the better sensitivity, which is expected as the optimal control gives the best possible performance, the general behavior of the optimal protocols is qualitatively similar to that of the reference protocols. In the short-$\tau$ regime which is relevant to high time resolution, we propose a detune protocol that only involves the smooth control field over the entire interrogation and performs comparably to the reference protocol. Given that discontinuities in the control field pose the primary experimental challenge, we expect the detune protocol to be practically useful. One bottleneck of meaningfully utilizing the time-resolution protocol is the need to generate the identical, yet unknown time-domain signal repeatably. We suggest that it can be used to calibrate the pulse distortion in fast flux control in superconducting qubits 10.1063/1.5133894; hellings2025calibratingmagneticfluxcontrol; Li_2025, which is an essential but cumbersome process for calibrating quantum gates, and is increasingly detrimental as the number of qubits gets larger. We discuss the potential to substantially reduce the total calibration time by leveraging our protocol to perform sequential measurements of the same distorted pulse. The rest of the paper is organized as follows. Section II defines the problem of time resolution and provides the necessary background. Section III presents our main results, including the general behavior from optimal control and a detailed description of the detune protocol. Section IV discusses practical aspects, including an outline of a realistic application. Section V provides a brief conclusion.
@@ II.1 Overview and problem statement
In this subsection we formulate the design of the time-resolution protocol for quantum sensing as an optimal control problem. The quantum sensor is modeled as a two-level system, which can physically correspond to an electron spin, an atom, or a superconducting qubit; it is described by the Hamiltonian PhysRevLett.133.210802
@@
where $\omega_{0}$ is the natural angular frequency of the quantum sensor, $\delta\omega$ is the field which we want to measure, and $u(t)$ is a scalar field which we can control. We shall take $\hbar\equiv 1$, $\omega_{0}\equiv 1$ in our simulations but explicitly keep $\omega_{0}$ in all expressions; within this convention energy is measured in units of $\hbar\omega_{0}$ and time $\omega_{0}^{-1}$. $|0\rangle$ and $|1\rangle$ are eigenstates of $\hat{\sigma}_{z}$ with eigenvalues of 1 and -1 respectively. They are eigenstates without external fields (i.e., $\delta\omega=u=0$) and will be referred to as the natural eigenstates of the quantum sensor.
The performance of the time-resolution protocol is quantified by its measurement sensitivity with respect to $\delta\omega$ given an (short) interrogation time $\tau$. Because any operation contributes to the interrogation time, an end-to-end description of the measurement process is required for performance evaluation. In particular the initial state and the basis of the final projection measurement have to be the sensor’s natural eigenstates. Without loss of generality we choose the initial state to be $|0\rangle$ and measure the probability that ends up with the same $|0\rangle$ state; the outcome probability can be expressed as
@@
$p_{0}$ is the probability with zero field (i.e., $\delta\omega=0$) and is independent of $\delta\omega$; the linear coefficient of $\delta\omega$, denoted as $\eta$, is the “measurement sensitivity” PhysRevLett.133.210802:
@@ II.2 Reference: resonant YX protocol
The reference protocol, which will be termed as the resonant YX protocol or simply the YX protocol, is the one proposed by Herb and Degen in Ref. PhysRevLett.133.210802; its control waveform is
@@
where $t_{\text{QSL}}=\frac{\pi}{u_{\text{max}}}$ is referred to as the quantum speed limit PhysRevLett.133.210802; QuantumSpeedLimit; Mandelstam_1945; Norma_1998; Deffner_2017. The rationale behind the protocol naming is as follows: the “resonant” indicates that the applied frequency is identical to the natural frequency; “YX” that in RWA, $\cos(\omega_{0}t+\frac{\pi}{2})\hat{\sigma}_{x}$ and $\cos(\omega_{0}t)\hat{\sigma}_{x}$ respectively lead to $\hat{\sigma}_{y}$ (Y operation) and $\hat{\sigma}_{x}$ (X operation) in the rotating frame. Overall the YX protocol identifies a critical interrogation time $t_{\text{QSL}}$, below which the protocol is composed of equal-duration Y and X operations, and above which the protocol involves a free evolution between Y and X operations.
@@
Here $K$ is a protocol dependent convolution kernel; for constant $\delta\omega$ one recovers $\eta=\int_{-\tau/2}^{\tau/2}\text{d}tK(t)$. In Appendix A (see also Supplementary of Ref. PhysRevLett.133.210802) we provide a simple numerical method to construct the kernel of Eq. (8). Once $K$ is known, the time resolution shorter than $\tau$ is in principle possible Herb_2025.
@@ II.3 Optimal control
In this subsection we summarize the important functions introduced in OCT. Numerous comprehensive reviews have been reported in the literature PRXQuantum.2.030203; PRXQuantum.2.010101, and its applications specific to the unitary qubit can be found in Ref. PhysRevA.111.042602. Here, special attention is devoted to the cost function that involves a derivative of the wave function.
@@
OCT is a powerful tool for dynamics-constrained optimization. Given the dynamics and the cost to minimize, it introduces the adjoint variables $|\pi(t)\rangle$, the control-Hamiltonian $\mathcal{H}_{\text{oc}}(t)$, and the switching function $\Phi(t)$ [defined shortly in Eqs. (12)] that are computationally accessible and accurately characterize the system response to some external or model variables. One non-trivial complication from the cost $\mathcal{C}_{\eta^{2}}$ is its dependence on $|\psi_{1}(\tau)\rangle=\frac{\partial}{\partial(\delta\omega)}|\psi_{0}(\tau)\rangle$. To cope with this additional dependence we augment the system by regarding $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$ as independent variables; more details can be found in Ref. PhysRevA.103.052607; PhysRevA.101.022320.
@@
We also consider maximizing QFI PhysRevA.103.052607 where the terminal cost is negative QFI:
@@
Eqs. (12) plus Eq. (13) [or Eq. (14)] completely determine quantities introduced by OCT. They are used to express two general optimality conditions control_type. First, the optimal control is
@@
We emphasize that although the physical system considered here is equivalent to a qubit with unitary dynamics [i.e., Eq. (1)], the augmented dynamics specified in Eq. (12c) is neither planar (i.e., two real-valued variables) nor unitary. For these reasons certain properties specific to the unitary qubit given in Ref. PhysRevA.111.042602 do not hold anymore. Particularly we point out that the BB control does not have identical bang duration and the singular control does not require $u^{\text{sing}}=0$.
@@ II.4 Practical use of OCT functions
We conclude this section by highlighting the usefulness of OCT. The most practically useful quantity is undoubtedly the switching function $\Phi$. By sequentially solving the dynamics twice, one forward in time for $|\psi\rangle$ and one backward $|\pi\rangle$, one is able to obtain the gradient of $\mathcal{C}$ with respect to all $u(t_{i})$ KHANEJA2005296. Three applications used in the work are highlighted. First $\Phi$ is primarily used to construct a numerical $u^{*}(t)$ by some gradient-based (or quasi-Newton) algorithm that involves
@@
For the free-form parametrization one divides $\tau$ into $N_{t}$ intervals and assumes a piecewise constant waveform. In our simulations $N_{t}$ ranges from 100 to 2000 (depending on $\tau$); we ensure that the result has a negligible change upon increasing $N_{t}$’s and that $N_{t}$ is sufficiently large so that all optimality conditions are numerically satisfied PhysRevA.103.052607. Empirically we find that when $u^{*}(t)$ contains both B and S controls, the convergence of S control is slow due to its vanishing gradient. In this case we can compute Eq. (16) at $n$th iteration, and replace the non-bang controls by the singular values:
@@
Two practical utilizations of control-Hamiltonian $\mathcal{H}_{\text{oc}}(t)\sim\frac{\partial\mathcal{C}}{\partial\tau}$ are outlined. First, the deviation from a constant $\mathcal{H}_{\text{oc}}(t)$ can be used to quantify the solution quality PhysRevA.103.052607. Second, in a damped quantum system $\mathcal{H}_{\text{oc}}=0$ offers a rigorous criterion for defining the optimal operation time for a quantum task PhysRevA.105.042621.